zahmel:

Which equation is equivalent to –k + 0.03 + 1.01k = –2.45 – 1.81k? –100k + 3 + 101k = –245 – 181k –100k + 300 + 101k = –245 – 181k –k + 3 + 101k = –245 – 181k –k + 300 + 101k = –245 – 181k

Tranquility:

Looking at the answer choices, it seems like they want to get rid of the decimals Multiply both sides by 100 100(-k + 0.03 + 1.01k) = 100*(-2.45 - 1.81k) Remember you have to distribute a(b+c) = (a*b) + (a*c)

zahmel:

100k+3+101k=-245-181k ?

Tranquility:

Almost, you forgot the - on the first 100k So it would be -100k + 3 + 101k = -245 - 181k

zahmel:

thank u

Tranquility:

No problem!